A 600-ton press used to produce composite material fuel cell components for automobiles using proton exchange membrane (PEM) technology can reduce the weight of enclosure parts up to 75%. At MARR = 12% per year, calculate (a) capital recovery and (b) annual revenue required. Installed cost = $3.8 million n = 12 years, Salvage value = $250,000, Annual operating costs _ $350,000 to start increasing by $25,000 per year


1
Expert's answer
2021-10-31T17:34:08-0400

Given;

Installed cost (P) =$3.8 million

n=12

Salvage value (S) = $250,000

Annual operating cost (A) = $350000

Operating cost I increases by(G) =$25000

MARR = 12%


a) The value of capital recovery

"CR=-P(A\/P,i,n)+S(A\/F,i,n)"

"=


CR= -3800000(A/P,12\%,12)+250000(A/F,12\%,12)" "=-\$3800000(0.16144)+\$250000(0.04144)"

"=-\$613472+\$10360=-\$603112"

The capital recovery is -$603112


b)Annual revenue required

"AW=CR+A" of "AOC"

"=-\$603112-\$350000-\$25000(A/G,12\%,12)"

"=-\$603112-\$350000-\$25000(4.1897)""=-\$603112-\$350000-\$104742.5 =-\$1057854.5"

The annual revenue required is -$1057854.5